OpenStudy (anonymous):

f(x)=sqrt2x-5 find f (7) A. 2 B. 3 C. -5 D. 7

OpenStudy (anonymous):

When they ask you to find f(7) and the equation is \[\large f(x)=\sqrt{2x-5}\] f(7) is the form of f(x), so what would the 7 be replacing in f(x)?

OpenStudy (anonymous):

@katlin95

OpenStudy (anonymous):

oh so 7 would be our answer right?

OpenStudy (anonymous):

No...I'm asking you aquestion unrelated to the answer. In maths, I never just go finding the answer in 2 seconds. You need stepping blocks before you can get to the answer...

OpenStudy (anonymous):

a question*

OpenStudy (anonymous):

yep the 7 would be replacingit

OpenStudy (anonymous):

What's "it"?

OpenStudy (anonymous):

Is the 7 replacing the moon or something. You're not being specific.

OpenStudy (anonymous):

f(x)

OpenStudy (anonymous):

f(x) and f(7)...what's the difference between those two figures. I'm just asking you to look at it as a picture and you're trying to spot the difference. what's different. Is the letter f different to each other, are the brackets different? What's different between f(x) and f(7)?

OpenStudy (anonymous):

That's a very straight forward question for children under 7. You should be able to get this in a flash.

OpenStudy (anonymous):

7 replaces x

OpenStudy (anonymous):

Excellent.

OpenStudy (anonymous):

So \[\large f(7)=\sqrt{2(7)-5}\]

OpenStudy (anonymous):

That's in the form of this \[\large f(x)=\sqrt{2x-5}\] Correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Everytime you see the term x, you sub in 7. So now could you just calculate the number that's inside the sqrt sign for me?

OpenStudy (anonymous):

i got 2x-5 still

OpenStudy (anonymous):

We're using this... -_-... \[\huge f(7)=\sqrt{2(7)-5}\] Does the question say find f(x)...Because if says that, then it's like child's play. You would just be rewriting what the question gave you...You must pay attention to detail and everything that's related to the QUESTION. The QUESTION in every single subject is like the holy bible in religious terms. The question is the key to everything you do to work out the answer.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct. Good job. Remember that important basic: read the question. People make mistakes because they're too lazy to read the question. It's that simple. People smart enough to just focus on the question and what it's asking you to do, would obviously be the one's getting top marks.

OpenStudy (anonymous):

yay!